# Traffic Co-Simulation Framework Empowered by Infrastructure Camera Sensing and Reinforcement Learning

### v1

@@ III-A Overview
The process starts with using a SUMO road network with traffic lights as Q-learning agents utilizing the sumo_rl Python library [45] which converts SUMO networks to RL environments based on the stable baselines format, allowing the use of standard RL algorithms that the library offers. Each traffic light is an independent agent taking the current phase, density of incoming lanes, and queue length as the state. Each agent calculates its reward from its incoming lanes based on one of four reward functions tested. The trained list of agents is saved to a Python pickle file to be used in the co-simulation.
@@
III-B Vehicle Detection and Aggregation using Virtual Infrastructure Cameras
@@
The CARLA-SUMO co-simulation Python file contains a bridge helper object which is used to get the corresponding locations between the two environments since they each operate with their separate coordinate systems. This function is used to match the CARLA traffic lights to SUMO junctions and to obtain the location of the corresponding stopping line for each traffic light. The yaw angle for the camera is then calculated using the camera position and the stopping line position by taking the inverse tangent of the ratio of the difference in $y$ to the difference in $x$ coordinates. Pitch angle is set to $-10\degree$ and roll angle is kept zero, ensuring that the concerning lane for each camera appears in the middle of the image frame. Using this method cameras can be automatically spawned on all traffic lights while ensuring correct placement and orientation.
@@
III-C Reinforcement Learning Agent Training
@@
where $\alpha$ is the learning rate, $\gamma$ is a discount factor that determines the importance of future rewards, $r$ is the reward after taking action $a$ from state $s$, and $s^{\prime}$ is the new state after the action. $max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ refers to the maximum Q-value for the next state over all possible actions $a^{\prime}$.
@@
Agent: Each traffic light is a reinforcement learning agent with its own state space, action space, learning rate, discount factor, and exploration strategy. While these variables can be made different for experimentation, in this work we initialize all agents with the same parameters. The action and state spaces, however, depend on the incoming lanes and number of phases.
@@
State space: The state of the $i$th agent can be represented as $s_{i}=[p,m,\mathbf{d},\mathbf{q}]$. It is an observation consisting of the current phase index $p$, a boolean variable $m$ indicating whether minimum green time has elapsed, density vector $\mathbf{d}$ of each incoming lane, and queue vector $\mathbf{q}$ of each incoming lane. The density is defined as the number of vehicles divided by the length of the lane, while the queue is the number of stopped vehicles divided by the number of vehicles that could fit in the lane. As there is no assumption of communication between agents, any information external to an agent is not available to it for decision making.
@@
Action space: The action is the index of the green phase to be activated at the next time step. Each agent can have a different action space corresponding to the number of green phases it has. For example, a dedicated left turn phase will be considered as an extra green phase in addition to the phase that allows both through and yielding left turn traffic. The environment controlling the simulation ensures that the appropriate yellow phase is inserted between different phases, and there is no implementation of an all red phase.
@@
Reward functions: The reward is an integral part of reinforcement learning and should represent the desired collective goal of the process while being local to each agent. It is calculated after a minimum delay following an action since the outcome of the action will not immediately be reflected in the traffic state. This delay is set to be the same as the minimum green time. There are four well-known reward functions provided in the sumo-rl library: difference of wait times, average speed, queued vehicles, and pressure. The waiting time reward is calculated as the sum of the waiting time of all vehicles in all incoming lanes in an intersection minus the same sum from the previous time the reward function was called. The average speed function is the time mean speed of all vehicles in incoming and outgoing lanes of the intersection normalized over the maximum allowed speed. The queue reward is the negative of the total vehicles queued at the intersection. Pressure reward is the difference between the number of vehicles leaving and approaching the intersection.
@@
III-D Reinforcement Learning Agent Deployment in SUMO
@@
III-E Synchronization and Component Integration
@@ III-E1 Time Synchronization
III-E2 Coordinate System Alignment
@@
III-E3 Camera Data as Real-Time Input to Traffic Light Agents
Virtual cameras are mounted on traffic light poles within the CARLA simulation. These cameras capture images at a frequency of once per second. The images are then processed using the YOLOv5 object detection algorithm to identify vehicles in the oncoming lanes. The detected vehicle information is aggregated and used to update the state of the reinforcement learning agents. During this computation the simulation does not advance and the data is synchronized to the simulation once all states are updated.
@@ III-E3 Camera Data as Real-Time Input to Traffic Light Agents
III-E4 Action Execution and Feedback Loop
@@
III-E5 Error Handling and Robustness
@@
IV Experimental Analysis
IV-A Platform and Parameter Settings
The experiments were executed on a desktop computer with 13th Gen Intel® Core™ i9-13900K CPU, 64 GB RAM, and NVIDIA RTX 4090 graphics card. The road network chosen for traffic signal optimization was ‘Town04’, a map included in the CARLA software. It contains a downtown grid-like region surrounded by a highway. Entry and exit from the downtown are mediated by traffic lights, and the whole map contains 12 traffic lights in total. As part of the RL process, each traffic light is a Q-learning agent with access to its local state observation of the current phase, minimum green time indication, and the density and queue for each incoming lane. We manually specify traffic flow demands to cause congestion in the downtown grid and ample activity into and out of that region. These flows are defined as origin and destination, so the routes are not fixed and the vehicles dynamically determine their path from estimated travel time within the simulation. With static signals, one run of simulation lasts 2770 seconds in simulation time.
The multi-agent RL was run with each reward function for 300 episodes with $\alpha=0.1,\ \gamma=0.99$ and an initial $\varepsilon$ value of 0.05. The learning progress of the agents is shown in Figure 3a which shows the average speed of vehicles in the network over the training run. Agents with average speed and pressure based reward showed significant network-wide improvement during training. Figure 3b depicts the average wait times at one chosen intersection which also shows major variation for pressure and average speed based reward, and only minor improvement for queue based reward. The process resulted in four sets of trained agents, each set corresponding to a reward type, which were then saved as Python pickle files to be later used for evaluation in the co-simulation.
@@ IV-A Platform and Parameter Settings
The synchronization file for CARLA-SUMO co-simulation was modified to carry out the experiment. A camera was spawned on each traffic light pole resulting in 43 total cameras, and the corresponding lane was matched using the stop location coordinates. Each camera updates an image in the form of a numpy array which is then used as input to a pytorch instance of YOLOv5 to detect vehicles within $240\times 480$ rectangle in the center of the frame. If the confidence score of the detection is higher than 0.5, it is counted as a vehicle belonging to the respective lane. Another function collects the required information and calculates density and queue length for each intersection to form the state vector to be used as input for the trained agents. If the agent Q-table does not contain the state, the action is not updated in the simulation and its phase does not change. Otherwise, the phase is set to the recommended one after going through the appropriate yellow phase.
There were two traffic demand scenarios tested in the co-simulation: medium and heavy. Both had the same routes but the heavy demand had higher vehicles per hour flow (591 vehicles at 200 to 400 veh/hr) for all routes as compared to the medium demand (387 vehicles at 100 to 300 veh/hr). The origin and destination for the demands were the same as the ones used for the MARL but differed in start and end times to ensure that we were not simply testing on the training scenario while also avoiding radically different traffic conditions. Congestion was observed in all scenarios, the last car entered at 1100 seconds into the simulations and the medium demand scenario lasted for 1700 seconds while heavy demand scenario ran for over 2800 seconds.
IV-B Co-Simulation Results
IV-B1 Traffic Signal Control with MARL
Baseline simulation with static timing traffic lights was carried out for the two demands and it showed an average vehicle speed of 6.11 m/s and a total travel time of 145,144 seconds in the medium traffic, and 4.57 m/s average speed with 355,699 seconds total travel time in the heavy traffic. Table II shows the performance of the trained RL agents controlling the traffic lights using counts from the vehicle detections to form the observation vector. Out of the four variants tested, the agents trained with the reward function based on the average speed of vehicles near the intersection performed the best in both test scenarios by every measure. It is important to mention that the rewards are not available at test time, only the observation of lane density and queue length is provided based on the camera sensing.
@@ IV-B1 Traffic Signal Control with MARL
Table III shows the results of the same simulations but with ground truth data of the lanes available to each agent. It can be seen that the correct knowledge of the state of the intersection results in a better performance from the trained agents except in one case. Wait time and pressure reward based agents take a significant performance hit when working with noisy sensing, while the effect on the agent trained with average speed reward performs less than 10% worse as compared to the ground truth scenario with medium traffic. The difference is more pronounced in heavy traffic. The relative robustness of agents with average speed reward indicates that while the other agents are learning first order relationships, speed based reward provides temporal insight to the Q-table. Since wait time and pressure are both directly linked to queue length and density, agents with perfect information develop an overfitted estimate of the Q-value of taking a certain action, whereas imperfect observations during testing cause a small difference in the state vector from one measurement to the next, resulting in a significantly different Q-value. On the other hand, being rewarded with average speed, an agent will prioritize staying in the same phase for longer if the state does not vary too much between observations.
Queue length based reward proved to be the worst performer in both cases but the drastic difference with ground truth information is definitely anomalous as the camera detection version had better metrics. A closer look at the simulation revealed that during congestion the traffic backs up onto the highway, which is not part of the short lane being observed by the traffic light. Therefore, the agent always observes that the queue is full, and any action it takes will not change the size of the queue at the next measurement. In other words, the expected reward is contained in the observation state itself. As a consequence, that particular agent never learns the real value of giving the green phase to the incoming traffic. Interestingly, the reason that the camera detection version performs better is that due to a failure to detect all the cars, or detecting extra cars in the adjacent lane, the count deviates from the ground truth causing different rows of the Q-table to be accessed, forcing the agent to change the phase.
TABLE II: Simulation performance measures of statically configured traffic lights as compared to agents trained using different reward functions using computer vision (YOLOv5) for observation
@@
TABLE III: Simulation performance using trained agents with ground truth observations
Wait time Average speed Queue Pressure
@@
Mean speed (m/s) 9.48 9.84 4.69 9.39
Total travel time (s) 79,451 76,316 243,401 80,846
Mean waiting time (s) 21.1 19.4 225 20.3
Mean time lost (s) 58.7 50.6 481 62.3
Mean trip duration (s) 205 197 629 209
@@
Mean speed (m/s) 7.00 7.11 3.67 5.68
Total travel time (s) 163,782 158,245 559,992 227,577
Mean waiting time (s) 42.3 36.3 400 108
Mean time lost (s) 136 127 804 243
Mean trip duration (s) 277 268 947 385
IV-B2 Video Detection

### v2

@@ 3.1 Overview
The process starts with using a SUMO road network with traffic lights as Q-learning agents utilizing the sumo_rl Python library (Alegre, 2019) which converts SUMO networks to RL environments based on the stable baselines format, allowing the use of standard RL algorithms that the library offers. Each traffic light is an independent agent taking the current phase, occupancy of incoming lanes, and queue length as the state. Each agent calculates its reward from its incoming lanes based on one of four reward functions tested. The trained list of agents is saved to a Python pickle file to be used in the co-simulation.
@@
3.2 Vehicle Detection and Aggregation using Virtual Infrastructure Cameras
@@
The CARLA-SUMO co-simulation Python file contains a bridge helper object which is used to get the corresponding locations between the two environments since they each operate with their separate coordinate systems. This function is used to match the CARLA traffic lights to SUMO junctions and to obtain the location of the corresponding stopping line for each traffic light. The yaw angle for the camera is then calculated using the camera position and the stopping line position by taking the inverse tangent of the ratio of the difference in $y$ to the difference in $x$ coordinates. Pitch angle is set to $10\degree$ and roll angle is kept zero, ensuring that the concerning lane for each camera appears in the middle of the image frame. Using this method cameras can be automatically spawned on all traffic lights while ensuring correct placement and orientation.
@@
3.3 Reinforcement Learning Agent Training
@@
where $\alpha$ is the learning rate, $\gamma$ is a discount factor that determines the importance of future rewards, $r$ is the reward after taking action $a$ from state $s$, and $s^{\prime}$ is the new state after the action. $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ refers to the maximum Q-value for the next state over all possible actions $a^{\prime}$.
@@
Agent: Let agent $i\in\{1,...,N\}$ control a signalized intersection. Each traffic light is a reinforcement learning agent with its own state space, action space, learning rate, discount factor, and exploration strategy. While these variables can be made different for experimentation, in this work we initialize all agents with the same parameters. The action and state spaces, however, depend on the incoming lanes and number of phases.
State space: The state of the $i$th agent at time $t$ can be represented as
$s_{i}^{t}=\left(p_{i}^{t},m_{i}^{t},\mathbf{d}_{i}^{t},\mathbf{q}_{i}^{t}\right)$
@@
It is an observation consisting of the current phase index $p^{t}_{i}\in\{0,1,...P-1\}$, a boolean variable $m^{i}_{t}\in\{0,1\}$ indicating whether minimum green time has elapsed, occupancy vector $\mathbf{d}^{t}_{i}\in\mathbb{N}^{L_{i}}$ of each incoming lane $L$ at intersection $i$, and queue vector $\mathbf{q}^{t}_{i}\in\mathbb{N}^{L_{i}}$ of each incoming lane. The occupancy is defined as the number of vehicles approaching the intersection from each incoming lane, while the queue is the number of stopped vehicles in each incoming lane. As there is no assumption of communication between agents, any information external to an agent is not available to it for decision making.
@@
Action space: The action is the integer index of the green phase to be activated at the next time step. Every agent can have a different action space corresponding to the number of green phases it has. Each agent $i$ selects an action $a^{t}_{i}\in\mathcal{A}_{i}$, where the action space is $\mathcal{A}_{i}=\left\{0,1,\ldots,G_{i}-1\right\}$. Here, $G_{i}$ is the number of valid green phases for intersection $i$. Each element corresponds to a unique phase (e.g., protected left turn, through traffic), as defined in the SUMO network. The environment controlling the simulation ensures that the appropriate yellow phase is inserted between different phases, and there is no implementation of an all red phase. For intersections with complex movements, each legally permissible combination of movements (e.g., a protected left turn) is defined as a unique phase with a corresponding index in SUMO rather than a sub-phase.
@@
Reward functions: The reward is an integral part of reinforcement learning and should represent the desired collective goal of the process while being local to each agent. It is calculated after a minimum fixed delay, $\Delta t$, following an action since the outcome of the action will not immediately be reflected in the traffic state. This delay is set to be the same as the minimum green time. Let $r_{i}^{t}$ denote the reward received by agent $i$ at time $t$. There are four well-known reward functions provided in the sumo-rl library and we test each one of them separately:
  • Difference of wait times calculated as the sum of the waiting time of all vehicles in all incoming lanes in an intersection minus the same sum from the previous time the reward function was called.
    $r_{i}^{t}=\sum_{\ell\in\mathcal{L}_{i}^{\text{in }}}W_{\ell}^{t}-\sum_{\ell\in\mathcal{L}_{i}^{\text{in }}}W_{\ell}^{t-\Delta t}$
    where $W^{t}_{\ell}$ is the total waiting time of vehicles on lane $\ell$ at time $t$.
  • Average speed is the time mean speed of all vehicles in incoming and outgoing lanes of the intersection normalized over the maximum allowed speed.
    $r_{i}^{t}=\frac{1}{\left|\mathcal{U}_{i}^{t}\right|}\sum_{u\in\mathcal{U}_{i}^{t}}\frac{v_{u}^{t}}{v_{\max}}$
    where $v^{t}_{u}$ is the speed of vehicle $u$, $v_{\text{max}}$ is the maximum allowed speed, and $\mathcal{U}_{i}^{t}$ is the set of vehicles in the vicinity of intersection $i$.
  • Queue length reward is the negative of the total vehicles queued at the intersection.
    $r_{i}^{t}=-\sum_{\ell\in\mathcal{L}_{i}^{\text{in }}}n_{\ell,\text{ stopped }}^{t}$
  • Pressure reward is the difference between the number of vehicles leaving and approaching the intersection.
    $r_{i}^{t}=\sum_{\ell\in\mathcal{L}_{i}^{\text{out }}}n_{\ell}^{t}-\sum_{\ell\in\mathcal{L}_{i}^{\text{in }}}n_{\ell}^{t}$
[…]
@@
3.4 Reinforcement Learning Agent Deployment in SUMO
@@
3.5 Synchronization and Component Integration
@@ 3.5.1 Time Synchronization
3.5.2 Coordinate System Alignment
@@
The module is used to access traffic light pole locations in the CARLA environment, spawning virtual cameras at these locations on a predefined height above the existing light pole. The IDs and coordinates of the stopping line of each lane are matched between CARLA and SUMO to give the appropriate label to the deployed camera. The camera yaw angle is calculated from the vector between the light pole and the stopping line to ensure it is properly aimed. This automated setup is a key technical contribution that makes scaling to larger networks feasible, a step beyond single-intersection or manually configured testbeds.
3.5.3 Camera Data as Real-Time Input to Traffic Light Agents
The virtual cameras were configured to capture images at a frequency of once per second. The images are then processed using the YOLOv5 object detection algorithm to identify vehicles in the oncoming lanes. During this computation the simulation does not advance and the data is synchronized to the simulation once all states are updated.
The MARL agents in our framework, trained using the sumo-rl library, require a structured state vector comprising features such as lane occupancy and queue length, typically derived from ground-truth simulator data. In contrast, the CARLA environment provides only raw image data from virtual cameras. Bridging this modality gap posed a key integration challenge. To address it, we developed a custom vision-to-state translation module that processes images from multiple intersection-mounted cameras using a YOLO-based object detector. The module filters detections by vehicle type, restricts analysis to a defined region of interest, and aggregates counts across cameras serving the same intersection. These counts are then mapped to the appropriate SUMO lane indices and transformed into occupancy and queue metrics compatible with the pre-trained agents’ Q-table. This translation layer is essential to enabling reinforcement learning control based on visual perception and represents a core technical contribution of our system.
@@ 3.5.3 Camera Data as Real-Time Input to Traffic Light Agents
3.5.4 Action Execution and Feedback Loop
@@
The co-simulation operates at 20 Hz to maintain smooth vehicle dynamics, but real-time object detection on 43 cameras is computationally infeasible. To address this, we implemented a control loop that periodically pauses the simulation (e.g., every 5 seconds) to perform batched detection across all cameras. Upon resuming, the system ensures data consistency, allowing agents to act on a coherent snapshot of the environment. This asynchronous coordination is essential for practical vision-in-the-loop control.
3.5.5 Error Handling and Robustness
@@
4 Experimental Analysis
4.1 Platform and Parameter Settings
The experiments were executed on a desktop computer with 13th Gen Intel® Core™ i9-13900K CPU, 64 GB RAM, and NVIDIA RTX 4090 graphics card. The road network chosen for traffic signal optimization was “Town04”, a map included in the CARLA software. It contains a downtown grid-like region surrounded by a highway. The East-West signal spacing is about 40-50 meters, while the North-South spacing is 50-70 meters. Entry and exit from the downtown are mediated by traffic lights, and the whole map contains 12 traffic lights in total. As part of the RL process, each traffic light is a Q-learning agent with access to its local state observation of the current phase, minimum green time indication, and the occupancy and queue for each incoming lane. The traffic signals operate independently: they are uncoordinated, possess no preprogrammed offsets, and do not exchange information with one another. The 9 downtown signals share the same cycle length in the case of static timing experiments. We manually specify traffic flow demands to cause congestion in the downtown grid and ample activity into and out of that region. These flows are defined as origin and destination, so the routes are not fixed and the vehicles dynamically determine their path from estimated travel time within the simulation. With static signals, one run of simulation lasts 2770 seconds in simulation time.
The learning progress of the agents is shown in Figure 3a which shows the average speed of vehicles in the network over the training run. Agents with average speed and pressure based reward showed significant network-wide improvement during training. Figure 3b depicts the average wait times at one chosen intersection which also shows major variation for pressure and average speed based reward, and only minor improvement for queue based reward. The process resulted in four sets of trained agents, each set corresponding to a reward type, which were then saved as Python pickle files to be later used for evaluation in the co-simulation.
4.1.1 Hyperparameter Tuning
To optimize the training process for the various MARL agents across different reward functions and environmental conditions, a systematic hyperparameter tuning approach was employed. Specifically, the learning rate ($\alpha$) was incrementally adjusted from 0.1 to 0.001, the discount factor ($\gamma$) was varied within the range of 0.5 to 0.99, and the initial exploration rate ($\varepsilon$) was tuned from 0 to 0.05. A consistent set of hyperparameters was maintained for all experiments, encompassing different reward functions and traffic flow scenarios, to ensure comparability of results. The number of training episodes was also varied up to 5000. Having found satisfactory values that give repeatable results, the multi-agent RL was ultimately run with each reward function for 300 episodes with $\alpha=0.0071,\ \gamma=0.97$ and an initial $\varepsilon$ value of 0.05 which decays to 0.005.
@@ 4.1.1 Hyperparameter Tuning
4.1.2 Co-simulation Testing Parameters
The synchronization file for CARLA-SUMO co-simulation was modified to carry out the experiment. A camera was spawned on each traffic light pole resulting in 43 total cameras, and the corresponding lane was matched using the stop location coordinates. Each camera updates an image in the form of a numpy array which is then used as input to a pytorch instance of YOLOv5 to detect vehicles within $240\times 480$ rectangle in the center of the frame. If the confidence score of the detection is higher than 0.5, it is counted as a vehicle belonging to the respective lane. Another function collects the required information and calculates occupancy and queue length for each intersection to form the state vector to be used as input for the trained agents. If the agent Q-table does not contain the state, the action is not updated in the simulation and its phase does not change. Otherwise, the phase is set to the recommended one after going through the appropriate yellow phase.
There were two traffic demand scenarios tested in the co-simulation: medium and heavy. Both had the same routes but the heavy demand had higher vehicles per hour flow (591 vehicles at 200 to 400 veh/hr) for all routes as compared to the medium demand (387 vehicles at 100 to 300 veh/hr). The origin and destination for the demands were the same as the ones used for the MARL but differed in start and end times to ensure that we were not simply testing on the training scenario while also avoiding radically different traffic conditions. Congestion was observed in all scenarios, the last car entered at 1100 seconds into the simulations and the medium demand scenario lasted for 1700 seconds while heavy demand scenario ran for over 2800 seconds. The most congested signal in the medium demand scenario experienced an average West-East demand ($v$) of 700 veh/hr while having a theoretical capacity ($C$) of 1800 veh/hr, and with a static green time ($g$) of 22 seconds out of a cycle length ($T$) of 50 seconds. The $v/C$ ratio can be calculated as
$\frac{v}{C}=\frac{v}{s\times g/T}=0.88$
which is a level of service (LOS) ‘E’, meaning unstable flow with high delays. Consistent queue spillback was observed for this intersection in the baseline simulation. Similarly, for the heavy demand scenario the same intersection received a demand of 1000 veh/hr, resulting in a $v/C$ ratio of 1.26, which is a considerably oversaturated LOS ‘F’.
4.2 Co-Simulation Results
4.2.1 Traffic Signal Control with MARL
Baseline simulation with static timing traffic lights was carried out for the two demands and it showed an average vehicle speed of 6.11 m/s and a total travel time of 145,144 seconds in the medium traffic, and 4.57 m/s average speed with 355,699 seconds total travel time in the heavy traffic. Table 2 shows the performance of the trained RL agents controlling the traffic lights using counts from the vehicle detections to form the observation vector. Out of the four variants tested, the agents trained with the reward function based on the average speed of vehicles near the intersection performed the best in both test scenarios by every measure. It is important to mention that the rewards are not available at test time, only the observation of lane occupancy and queue length is provided based on the camera sensing.
@@ 4.2.1 Traffic Signal Control with MARL
Table 3 shows the results of the same simulations but with ground truth data of the lanes available to each agent. This includes an adaptive actuated traffic light controller that works based on the time-gap between vehicles. It operates by prolonging the green phase for consecutive traffic, with a maximum duration of 60 seconds. In both traffic scenarios the best trained agents were able to outperform the actuated signals, except in the metric of average speed during heavy traffic. This can be attributed to the actuated signal’s tendency to prioritize overall vehicle movement by extending green phases for major thoroughfares, which can lead to a higher average speed across the network. However, this often comes at the cost of longer waiting times for vehicles on minor flows, as the system might not be as effective at minimizing individual vehicle delays at intersections. In contrast, the MARL agent, even when trained with an average speed reward, appears to learn a more sophisticated policy that balances network throughput with minimizing individual vehicle delays. For heavy traffic, minimizing stopped time (waiting time) is paramount for preventing gridlock and improving overall network efficiency, even if it means slightly lower peak speeds. This highlights a fundamental trade-off in how different adaptive control strategies optimize traffic, and suggests that reinforcement learning can offer distinct advantages in prioritizing local congestion relief and driver waiting times in extreme congestion.
It can be seen from Table 3 that the correct knowledge of the state of the intersection results in a better performance as compared to the same agents in Table 2 except in one case. Wait time and pressure reward based agents take a significant performance hit when working with noisy sensing, while the effect on the agent trained with average speed reward performs less than 10% worse as compared to the ground truth scenario with medium traffic. The difference is more pronounced in heavy traffic. The relative robustness of agents with average speed reward indicates that while the other agents are learning first order relationships, speed based reward provides temporal insight to the Q-table. Since wait time and pressure are both directly linked to queue length and occupancy, agents with perfect information develop an overfitted estimate of the Q-value of taking a certain action, whereas imperfect observations during testing cause a small difference in the state vector from one measurement to the next, resulting in a significantly different Q-value. On the other hand, being rewarded with average speed, an agent will prioritize staying in the same phase for longer if the state does not vary too much between observations.
Table 2: Simulation performance measures of statically configured traffic lights as compared to agents trained using different reward functions using computer vision (YOLOv5) for observation
@@
Table 3: Simulation performance using actuated lights and trained agents agents trained using different reward functions with ground truth observations
Actuated Wait time Average speed Queue Pressure
@@
Mean speed (m/s) 9.37 9.48 9.84 4.69 9.39
Total travel time (s) 85,273 79,451 76,316 243,401 80,846
Mean waiting time (s) 23.48 21.1 19.4 225 20.3
Mean time lost (s) 73.3 58.7 50.6 481 62.3
Mean trip duration (s) 220 205 197 629 209
@@
Mean speed (m/s) 8.03 7.00 7.11 3.67 5.68
Total travel time (s) 181,715 163,782 158,245 559,992 227,577
Mean waiting time (s) 92.96 42.3 36.3 400 108
Mean time lost (s) 163 136 127 804 243
Mean trip duration (s) 308 277 268 947 385
Queue length based reward proved to be the worst performer in both cases but the drastic difference with ground truth information is definitely anomalous as the camera detection version had better metrics. A closer look at the simulation revealed that during congestion the traffic backs up onto the highway, which is not part of the short lane being observed by the traffic light. Therefore, the agent always observes that the queue is full, and any action it takes will not change the size of the queue at the next measurement. As a consequence, that particular agent never learns the real value of giving the green phase to the incoming traffic. Interestingly, the reason that the camera detection version performs better is that due to a failure to detect all the cars, or detecting extra cars in the adjacent lane, the count deviates from the ground truth causing different rows of the Q-table to be accessed, forcing the agent to change the phase.
4.2.2 Video Detection